OpenStudy (anonymous):

What is the area of an equilateral triangle that has sides of 2 yards and a height of 4.25 feet?

OpenStudy (nurali):

First convert yards to feet: 3ft/yd --> 2yd = 6 ft Area triangle = 1/2*b*h = 1/2*6*4.25 = 12.75 ft.2 or 4.25 yd2

OpenStudy (anonymous):

Area of an equilateral triangle\[\sqrt{3}/4 \times side ^{2} \]